OpenStudy (anonymous):

Just want help doing it! Don't want you to give me the answer. What are the solutions of 3x2 + 4x = –5?

OpenStudy (anonymous):

is 3x2 3 times 2?

OpenStudy (anonymous):

or 3x (times) 2?

OpenStudy (jim766):

add 5 to both sides to make it equal to zero tell me what you get

OpenStudy (anonymous):

sorry 3x^2

OpenStudy (anonymous):

ohhh :O

OpenStudy (anonymous):

You know this is your 5th question like that . Every now and then, its good to try something yourself

OpenStudy (anonymous):

3x^2+4x+5=0 @Jim766 @Luis_Rivera I have done alot myself. I am just asking for help to understand scene I have to learn this all on my own with out a teacher if you got a problem with me asking questions on getting help then don't look at my stuff!

jimthompson5910 (jim_thompson5910):

You can't factor, so you'll have to use the quadratic formula \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] in this case a = 3 b = 4 c = 5

OpenStudy (anonymous):

in problems like this just set the function to 0 and use quadratic formula

OpenStudy (jim766):

the answer will not be real....b^2 - 4ac is negative are you sure you copied the problem correctly?

OpenStudy (anonymous):

@jim_thompson5910 & @galacticwavesXX well I'm stuck here I crossed out two answers but i got two left but I don't understand. \[\frac{-4 \pm \sqrt{4^2-4(3)(5)}} {2(3)}\] = \[\frac{-4 \pm \sqrt{16-60}} {6}\] =\[\frac{-4 \pm \sqrt{44}} {6}\] and this is where i'm stuck. Answer are- A.\[\frac{-4 \pm i\sqrt{66}} {6}\] B. \[\frac{-2 \pm 2i\sqrt{66}} {3}\] c.\[\frac{-2 \pm i\sqrt{11}} {3}\] D. \[\frac{-4 \pm i\sqrt{11}} {6}\] What do i do from here?

jimthompson5910 (jim_thompson5910):

You should be getting this \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(4)\pm\sqrt{(4)^2-4(3)(5)}}{2(3)}\] \[\Large x = \frac{-4\pm\sqrt{16-(60)}}{6}\] \[\Large x = \frac{-4\pm\sqrt{-44}}{6}\] Does that help?

jimthompson5910 (jim_thompson5910):

That's not the final answer, but it will eventually get you there

OpenStudy (anonymous):

how did you get -44?

jimthompson5910 (jim_thompson5910):

16 - 60 = -44

OpenStudy (anonymous):

yeah nevermind I subtracted wrong lol. But how do I go from here?

jimthompson5910 (jim_thompson5910):

well \[\Large \sqrt{-44} = \sqrt{-1*44}\] \[\Large \sqrt{-44} = \sqrt{-1*4*11}\] \[\Large \sqrt{-44} = \sqrt{-1}*\sqrt{4}*\sqrt{11}\] \[\Large \sqrt{-44} = i*2*\sqrt{11}\] \[\Large \sqrt{-44} = 2i*\sqrt{11}\]

jimthompson5910 (jim_thompson5910):

So \[\Large x = \frac{-4\pm\sqrt{-44}}{6}\] turns into \[\Large x = \frac{-4 \pm 2i*\sqrt{11}}{6}\]

OpenStudy (anonymous):

So then would the answer be C?

OpenStudy (jim766):

yes

jimthompson5910 (jim_thompson5910):

correct, it's C

OpenStudy (anonymous):

Okay thank you :)

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

3x2+4x+5=0 X=(-b±√(b^2-4ac))/2a X=(-4±√(4^2-4*3*5))/(2*3) X=(-4±√(16-60))/6 X=(-4±√(-44))/6 X=(-4±2×11i)/6 X=(-4)/6±22i/6 X=(-2)/3±11i/3